# Requirements Coverage-Guided Minimization for Natural Language Test Cases

### v2

@@ 8. Conclusion
In this paper, we presented RTM, a novel and practical approach for test suite minimization guided by requirements coverage. As is often the case in domains where functional safety is crucial and must be demonstrated through adherence to standards, test cases are specified in natural language (test steps), traced to requirements, and implemented manually.
RTM combines the text embedding technique, distance function, and a GA to effectively reduce test suite size while strictly preserving full requirement coverage under a fixed budget constraint. We investigate three preprocessing methods for test cases, seven text embedding techniques, three distance functions, and three initialization strategies for the GA. Its evaluation on an industrial automotive dataset demonstrates its superior performance in Fault Detection Rate (FDR) across various minimization budgets. Compared with multiple baselines, while being efficient in runtime, RTM achieves higher FDR across most minimization budgets, making it a better choice in many practical contexts. Furthermore, our exploration of test suite redundancy levels revealed that higher redundancy significantly improves FDR, while higher redundancy combined with low minimization budgets tends to significantly increase the risk of undetected faults, highlighting the importance of considering redundancy characteristics in test suite minimization.

### v3

@@ 8. Conclusion
In this paper, we presented RTM, a novel and practical approach for TSM guided by requirements coverage. As is often the case in domains where functional safety is crucial and must be demonstrated through adherence to standards, test cases are specified in natural language (test steps), traced to requirements, and implemented manually.
RTM combines the text embedding technique, distance function, and a GA to effectively reduce test suite size while strictly preserving full requirement coverage under a fixed budget constraint. We investigate three preprocessing methods for test cases, seven text embedding techniques, three distance functions, and three initialization strategies for the GA. Its evaluation on an industrial automotive dataset demonstrates its superior performance in FDR across various minimization budgets. Compared with multiple baselines, while being efficient in runtime, RTM achieves higher FDR across most minimization budgets, making it a better choice in many practical contexts. Furthermore, our exploration of test suite redundancy levels revealed that higher redundancy significantly improves FDR, while higher redundancy combined with low minimization budgets tends to significantly increase the risk of undetected faults, highlighting the importance of considering redundancy characteristics in TSM.